# ChartQA-X: Generating Explanations for Visual Chart Reasoning

### v2

@@ 3.4 Verifying Correctness
To verify the correctness of top-scoring explanations, we use the five models that did not generate the explanation. Each model independently predicts whether the explanation is correct or incorrect. If the majority label it as correct, the explanation is included in the ChartQA-X dataset. Otherwise, we evaluate the next-highest ROSCOE-scoring explanation using the same five models. This process repeats until an explanation is approved by the majority. If none are accepted, we regenerate explanations using all six models and restart the evaluation. Figure 4(a) shows the contribution of each model to the ChartQA-X dataset, while Figure 4(b) presents the word count distribution of the selected explanations.
@@ 5.1 Baselines and Metrics
We evaluate six open-source models, including LLaVA 1.6 [23], Qwen2-VL [40], InternVL-2.5 [6], Deepseek-VL [24], Phi-3 [1], and CogVLM [41], and a proprietary model, GPT-4o [15], on the ChartQA-X test set. We also fine-tune LLaVA 1.6, Qwen2-VL, and InternVL-2.5 on the ChartQA-X training set to establish three baseline models. All experiments are conducted with a batch size of 1 on 4 NVIDIA A100 (80GB) GPUs. Fine-tuning is performed on 8 NVIDIA H100 (80GB) GPUs and takes approximately six hours. Inference time on the benchmark is about one hour for all models, except CogVLM, which requires roughly 2.5 hours. For consistency, we use 7B-parameter versions of all models, except InternVL, which has 8B. We assess model performance using three criteria: (1) accuracy on the question-answering task, (2) explanation quality based on ROSCOE and n-gram metrics, including BLEU-4 [30], METEOR [3], ROUGE-L [22], and CIDEr [39], and (3) generalizability to unseen benchmarks after fine-tuning on ChartQA-X.
@@ 5.2 Results
Table 5 demonstrates the impact of fine-tuning on ChartQA-X across various chart and question types. Fine-tuned models significantly outperform their base counterparts. InternVL-2.5 fine-tuned on ChartQA-X achieves the highest overall accuracy of 79.78%. Substantial gains are observed across all chart types, with improvements of up to 18.96 percentage points (Qwen2-VL on line charts). On the reasoning subset, which is typically more challenging, ChartQA-X fine-tuning leads to notable boosts. For instance, Qwen2-VL improves from 34.33% to 48.01%. Even models with strong initial performance (e.g., InternVL-2.5) benefit, indicating the broad effectiveness of ChartQA-X. These results highlight ChartQA-X’s value in enhancing both factual and inferential understanding of visual charts. While prior papers report higher accuracy on the original ChartQA benchmark, our evaluation shows lower performance due to a key difference in our setup: we do not provide the underlying data tables as input, which makes the task significantly more challenging.
Table 4 shows that fine-tuning significantly improves explanation quality on the ChartQA-X test set, as measured by n-gram metrics. For example, Qwen2-VL shows significant absolute gains on CIDEr (+24.57), BLEU-4 (+14.01), and METEOR (+9.13), while InternVL-2.5 improves by +20.78, +10.76, and +4.75, respectively. Table 7 in Appendix A.3 further shows that fine-tuned models achieve significantly higher ROSCOE scores on the ChartQA-X test set.
@@ 6 Conclusion
We introduce ChartQA-X, the largest dataset of visual chart questions paired with detailed explanations. Alongside this dataset, we propose a novel explanation generation approach that leverages multiple models and evaluates outputs using ROSCOE metrics. Human evaluation indicates that ChartQA-X explanations are rated comparable to or better than human-written explanations across key dimensions of accuracy, clarity, logic, and overall quality. Furthermore, models fine-tuned on ChartQA-X significantly outperform existing baselines in generating chart explanations. These models also achieve higher accuracy in answering questions on external chart datasets, demonstrating strong generalization and robustness.
@@ A.2 Explanation Lengths
A.3 ROSCOE Scores
@@
Tables 7, 8, 9, and 10 present ROSCOE scores for explanations generated by various VLMs on ChartQA-X, DVQA, PlotQA, and FigureQA, respectively.
A.4 Qualitative Analysis
@@
Models FS FT IS IC SRC SFC PS PC GS AS State-of-the-Art Models Phi-3 [1] 0.865 0.939 0.873 0.925 0.678 0.792 0.995 0.947 0.946 0.885 CogVLM [41] 0.873 0.938 0.887 0.919 0.471 0.714 0.997 0.997 0.937 0.934 Deepseek-VL [24] 0.871 0.939 0.884 0.921 0.311 0.722 0.996 0.928 0.928 0.836 GPT-4o [15] 0.882 0.941 0.892 0.916 0.803 0.850 0.972 0.972 0.937 0.905 Models Fine-tuned on ChartQA-X InternVL-2.5 [7] 0.862 0.939 0.868 0.918 0.715 0.722 0.995 0.959 0.951 0.891      + ChartQA-X 0.894 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.032}$ 0.947 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.008}$ 0.877 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.009}$ 0.922 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.004}$ 0.797 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.082}$ 0.853 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.131}$ 0.994 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}-0.001}$ 0.963 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.004}$ 0.958 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.007}$ 0.904 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.013}$ LLaVA 1.6 [23] 0.884 0.945 0.877 0.901 0.515 0.858 0.989 0.991 0.941 0.873      + ChartQA-X 0.928 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.044}$ 0.949 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.004}$ 0.894 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.017}$ 0.918 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.017}$ 0.609 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.094}$ 0.920 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.062}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.008}$ 0.996 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.005}$ 0.980 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.039}$ 0.896 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.023}$ Qwen2-VL [40] 0.868 0.941 0.868 0.918 0.674 0.871 0.983 0.995 0.964 0.894      + ChartQA-X 0.907 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.039}$ 0.953 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.012}$ 0.884 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.016}$ 0.963 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.045}$ 0.902 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.228}$ 0.918 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.047}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.014}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.002}$ 0.970 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.006}$ 0.943 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.049}$
Table 7: ROSCOE scores on the ChartQA-X test set. Best scores are in bold, and second-best scores are underlined. FS: Faithfulness Step, FT: Faithfulness Token, IS: Informativeness Step, IC: Informativeness Chain, SRC: Source-Consistency, SFC: Self-Consistency, PS: Perplexity Step, PC: Perplexity Chain, GS: Grammar Step, and AS: Aggregate Score.

### v4

@@ 3.4 Verifying Correctness
To verify the correctness of the top-scoring explanations, we leverage the same six state-of-the-art vision-language models: LLaVA 1.6, Phi-3, CogVLM, Deepseek-VL, Qwen2-VL, and GPT-4o. For each explanation generated by a given model, the remaining five models serve as independent verifiers. Each verifier receives the input chart, question, ground-truth answer, data table, and candidate explanation, and produces a binary judgment, Correct or Incorrect. An explanation is accepted into the ChartQA-X dataset if a majority (i.e., at least three out of five) of the verifying models label it as correct. If the explanation fails to reach majority approval, we move to the next highest-scoring explanation based on its ROSCOE score and repeat the same verification process. This iterative evaluation continues until an explanation is accepted. If none of the initial explanations are accepted, we regenerate the candidate explanations from all six models and restart the process. Figure 4(a) shows the contribution of each model to the ChartQA-X dataset, while Figure 4(b) presents the word count distribution of the accepted explanations. See Appendix A.2 for the distribution of explanation word counts across the six VLMs.
@@ 5.1 Baselines and Metrics
We evaluate six open-source models, including LLaVA 1.6 [26], Qwen2-VL [43], InternVL-2.5 [6], Deepseek-VL [27], Phi-3 [1], and CogVLM [44], and a proprietary model, GPT-4o [16], on the ChartQA-X test set. We also fine-tune LLaVA 1.6, Qwen2-VL, and InternVL-2.5 on the ChartQA-X training set to jointly generate both the answer and an accompanying explanation in a single output sequence, establishing three baseline models. Each training instance is formatted as: “Answer: <answer>. Explanation: <explanation>”. The model is trained using a standard causal language modeling (LM) loss applied uniformly across the entire output sequence, without assigning distinct weights to the answer and explanation components. This joint generation approach promotes logical consistency between the answer and its explanation, enhancing interpretability, and reducing the risk of error propagation commonly associated with two-stage (answer-then-explanation) methods. At inference time, the model is prompted with:
[…]
This prompt is followed by the chart image and the question. All experiments are conducted with a batch size of 1 on 4 NVIDIA A100 (80GB) GPUs. Fine-tuning is performed on 8 NVIDIA H100 (80GB) GPUs and takes approximately six hours. Inference time on the test set is about one hour for all models, except CogVLM, which requires roughly 2.5 hours. For consistency, we use 7B-parameter versions of all models, except InternVL, which has 8B. We assess model performance using three criteria: (1) accuracy on the question-answering task, (2) explanation quality based on ROSCOE and n-gram metrics, including BLEU-4 [33], METEOR [3], ROUGE-L [25], and CIDEr [42], and (3) generalizability to unseen benchmarks after fine-tuning on ChartQA-X.
@@ 5.2 Results
Table 3 demonstrates the impact of fine-tuning on ChartQA-X across various chart and question types without data table in input. Fine-tuned models significantly outperform their base counterparts. InternVL-2.5 fine-tuned on ChartQA-X achieves the highest overall accuracy of 79.78%. Substantial gains are observed across all chart types, with improvements of up to 18.96 percentage points (Qwen2-VL on line charts). On the reasoning subset, which is typically more challenging, ChartQA-X fine-tuning leads to notable boosts. For instance, Qwen2-VL improves from 34.33% to 48.01%. Even models with strong initial performance (e.g., InternVL-2.5) benefit, indicating the broad effectiveness of ChartQA-X. These results highlight ChartQA-X’s value in enhancing both factual and inferential understanding of visual charts. While prior papers report higher accuracy on the original ChartQA benchmark, our evaluation shows lower performance due to a key difference in our setup: we do not provide the underlying data tables as input, which makes the task significantly more challenging. As shown in Table 4, incorporating the data table in the input substantially improves performance, with Qwen2-VL and LLaVA-1.6 achieving nearly twice their previous accuracy.
Table 5 further shows that fine-tuned models achieve significantly higher ROSCOE scores on the ChartQA-X test set. Table 6 shows that fine-tuning significantly improves explanation quality on the ChartQA-X test set, as measured by n-gram metrics. For example, Qwen2-VL shows significant absolute gains on CIDEr (+24.57), BLEU-4 (+14.01), and METEOR (+9.13), while InternVL-2.5 improves by +20.78, +10.76, and +4.75, respectively.
@@ 6 Conclusion
We introduce ChartQA-X, the largest dataset of visual chart questions paired with detailed explanations. Alongside this dataset, we propose a novel carefully-designed explanation generation pipeline that leverages multiple models and evaluates outputs using ROSCOE metrics. Human evaluation indicates that ChartQA-X explanations are rated comparable to or better than human-written explanations across key dimensions of accuracy, clarity, logic, and overall quality. Furthermore, models fine-tuned on ChartQA-X significantly outperform existing baselines in generating chart explanations. These models also achieve higher accuracy in answering questions on external chart datasets, demonstrating strong generalization and robustness.
Acknowledgments
This research was supported by the National Eye Institute (NEI) of the National Institutes of Health (NIH) under award number R01EY034562. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH. We thank ASU graduate students for contributing human-written chart explanations and ASU Research Computing (RC) for providing computing resources.
@@ A.2 Explanation Lengths
A.3 Sensitivity Analysis on ROSCOE Weights
To evaluate the robustness of our explanation selection method to different ROSCOE weight configurations, we conduct a sensitivity analysis using eight distinct weighting schemes. These include: (1) our original hand-tuned weights; (2) uniform weights across all nine metrics; (3) random weights sampled from a uniform distribution; (4) random weights from a Gaussian distribution; (5) fluency-focused weights with 0.08 for faithfulness step, 0.07 for faithfulness token, 0.1 for informativeness step and chain, self-consistency and source-consistency, and 0.15 for perplexity step and chain and grammar step; (6) faithfulness-focused weights with 0.30 for faithfulness step, 0.20 for faithfulness token, 0.15 for informativeness step and chain, 0.05 for source-consistency and self-consistency, 0.04 for perplexity step and chain and 0.02 for grammar step; (7) informativeness-focused weights with 0.10 for faithfulness step and token, 0.25 for informativeness step and chain, 0.15 for source-consistency, 0.05 for self-consistency, 0.025 for perplexity step and chain, and 0.05 for grammar step; and (8) a perturbed version of our original weights with small noise added with 0.18 for faithfulness step, 0.17 for faithfulness token, 0.14 for informativeness step, 0.16 for informativeness chain, 0.09 for source-consistency, 0.10 for self-consistency, 0.06 for perplexity step and 0.05 for perplexity chain, and 0.05 for grammar step.
Each weighting scheme is applied separately to compute aggregate ROSCOE scores and select the best explanation among the outputs of all six models for the entire dataset. We then compute the percentage of overlapping best explanation selections between all pairs of weighting sets. The results are shown in Table 9. The first row compares each alternative configuration with our proposed weights.
Compared Sets Set 2 Set 3 Set 4 Set 5 Set 6 Set 7 Set 8 Set 1 (Ours) 86.87 78.55 74.90 82.21 86.39 88.55 98.76 Set 2 82.17 76.10 94.30 74.86 81.49 87.03 Set 3 74.62 82.29 77.51 74.74 78.67 Set 4 74.94 71.08 67.83 75.66 Set 5 79.52 78.92 82.89 Set 6 79.70 84.79 Set 7 87.83
Table 9: Agreement (%) in explanation selection between different ROSCOE weight configurations. Each value indicates the percentage of matching explanation selections. The weight sets are described in Section A.3.
The analysis reveals that explanation selections are highly consistent across different weight sets. Agreement with our original configuration (Set 1) ranges from 74.90% (Set 4, random Gaussian) to 98.76% (Set 8, perturbed). Even randomly assigned weights (Sets 3 and 4) maintain agreement above 74%. Configurations focused on fluency, faithfulness, and informativeness (Sets 5–7) also perform similarly, with agreement rates over 82%. We also found that in 54.74% of the samples, all eight configurations selected the same explanations. These findings indicate that explanation selection using weighted ROSCOE is stable and not overly sensitive to weight configuration.
Models FS FT IS IC SRC SFC PS PC GS AS State-of-the-Art Models Phi-3 [1] 0.859 0.939 0.870 0.926 0.684 0.786 0.995 0.943 0.946 0.873 CogVLM [44] 0.877 0.939 0.882 0.919 0.659 0.700 0.990 0.939 0.909 0.864 Deepseek-VL [27] 0.876 0.940 0.890 0.920 0.679 0.710 0.994 0.933 0.891 0.867 GPT-4o [16] 0.854 0.937 0.861 0.915 0.728 0.847 0.994 0.964 0.954 0.881 Models Fine-tuned on ChartQA-X InternVL-2.5 [7] 0.864 0.939 0.869 0.915 0.748 0.733 0.994 0.937 0.945 0.873     + ChartQA-X 0.882 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.018}$ 0.940 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.001}$ 0.895 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.026}$ 0.926 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.011}$ 0.800 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.052}$ 0.839 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.106}$ 0.995 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.001}$ 0.966 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.029}$ 0.949 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.004}$ 0.900 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.013}$ LLaVA 1.6 [26] 0.889 0.949 0.880 0.922 0.529 0.847 0.989 0.950 0.931 0.872     + ChartQA-X 0.925 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.036}$ 0.952 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.003}$ 0.895 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.015}$ 0.930 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.008}$ 0.728 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.199}$ 0.984 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.137}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.008}$ 0.977 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.027}$ 0.955 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.024}$ 0.919 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.023}$ Qwen2-VL [43] 0.877 0.943 0.876 0.922 0.750 0.879 0.995 0.964 0.946 0.895     + ChartQA-X 0.905 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.028}$ 0.953 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.010}$ 0.892 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.016}$ 0.963 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.041}$ 0.901 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.151}$ 0.917 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.039}$ 0.997 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.002}$ 0.988 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.024}$ 0.957 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.011}$ 0.931 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}+0.049}$
Table 10: ROSCOE scores on the ChartQA-X test set with data table included as input. Best scores are in bold, and second-best scores are underlined. FS: Faithfulness Step, FT: Faithfulness Token, IS: Informativeness Step, IC: Informativeness Chain, SRC: Source-Consistency, SFC: Self-Consistency, PS: Perplexity Step, PC: Perplexity Chain, GS: Grammar Step, and AS: Aggregate Score.
A.4 ROSCOE Scores
@@
Tables 10, 11, 12, and 13 present ROSCOE scores for explanations generated by various VLMs on ChartQA-X with data table, DVQA, PlotQA, and FigureQA, respectively.
A.5 Qualitative Analysis